# A Calculus of Inheritance

### v4

@@ 4. Embedding the $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to inheritance-calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application:
@@
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
@@
$x$ ($\lambda$-bound, index $n$)   $\longrightarrow$ $\uparrow^{\!n}.\mathrm{argument}$
@@
$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$   $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$V_{1}\;V_{2}$ (tail call)   $\longrightarrow$ $\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
@@
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the tail-call rule, the fresh label $\mathrm{tail}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In both cases the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
An elementary semantics of the $\lambda$-calculus
Composing the translation $\mathcal{T}$ with the semantic functions of Section 3 induces a semantics on the $\lambda$-calculus itself. Given a $\lambda$-term $M$, one translates it to $\mathcal{T}(M)$ and then determines convergence behavior by querying $\ell\in\mathtt{properties}(p)$ via equations (1)–(6), following the $\mathrm{result}$ chain until the abstraction shape appears or the fixed point yields the empty set.
This induced semantics is more elementary than traditional approaches in the mathematical machinery it requires:
No function spaces:
Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
No substitution:
Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
No reduction rules:
There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975). The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required. Section 4.1 shows that the convergence behavior of the translated mixin tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract with respect to $\lambda$-definable contexts, meaning two $\lambda$-terms are equivalent under the translation iff they have the same Böhm tree.
Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics (w.r.t. $\lambda$-definable contexts) that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity. The mixin tree semantics achieves full abstraction with strictly simpler foundations than either.
@@ 4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
@@
in the least fixed point of Section 3.
@@
Theorem 4.3 (Full abstraction w.r.t. $\lambda$-definable contexts).
@@
Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2). Together, Adequacy and Full Abstraction establish that the translation $\mathcal{T}$ is a fully faithful embedding of the lazy $\lambda$-calculus into inheritance-calculus with respect to Böhm tree equivalence (Barendregt, 1984), using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: inheritance-calculus contexts can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathrm{inh}}|_{\mathrm{im}(\mathcal{T})})}$. This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem A.3).
@@ 4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in inheritance-calculus by inheriting new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3). The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
@@ 5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
@@
Following the Church-encoding convention (Yang, 2026), we define a factory containing the abstract product type and its two constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
@@
Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
To test that $2+3=5$, we needed equality on Nat values. Equality requires case analysis on the constructor; we use the Visitor pattern. $\mathrm{NatVisitor}$ adds a $\mathrm{Visitor}$ operation to each constructor, with a $\mathrm{Visit}$ field that delegates to the appropriate $\mathrm{VisitZero}$ or $\mathrm{VisitSuccessor}$ branch:
@@
$\quad\mathrm{Product}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{Visit}\mapsto\{\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitZero}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitZero}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitSuccessor}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitSuccessor}\}\}\}\}\}$
@@
$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Visitor}$:
@@
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\},$
@@
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\}\}\}$
@@
Each operation is an independent mixin that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited mixins. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential. In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix B), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
@@
A Church-encoded Nat is a record whose structure mirrors the constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:

### v6

@@ 4. Embedding the $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to inheritance-calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application. We extend the standard ANF grammar with a value binding $\mathbf{let}\;x=V\;\mathbf{in}\;M$, which gives a name to a value that is not an application result: [^5] Standard ANF allows values inline as operands; the value-binding production names them so that every subexpression of the translated program is accessible via a lexical reference (Appendix C).
@@
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results (and, optionally, naming values), and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
@@
$x$ ($\lambda$-bound, index $n^{\prime}$)    $\longrightarrow$ $\uparrow^{\!n^{\prime}}.\mathrm{argument}$
@@
$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$    $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},$
   $\longrightarrow$ $\mathrm{result}\mapsto\mathcal{T}(M)\}$
$\mathbf{let}\;x=V\;\mathbf{in}\;M$    $\longrightarrow$ $\{x\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$V_{1}\;V_{2}$ (tail call)    $\longrightarrow$ $\{\mathrm{tailCall}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},$
   $\longrightarrow$ $\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$
@@
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the application $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the value $\mathbf{let}$-binding rule, $x$ wraps $\mathcal{T}(V)$ under a $\mathrm{result}$ child so that the uniform accessor $x.\mathrm{result}$ yields $\mathcal{T}(V)$, the translated value itself. This extra wrapper is semantically inert but proof-relevant: it lets the same $\mathrm{result}$-following argument used for applications extend uniformly to value $\mathbf{let}$-bindings in the adequacy proof below. In the tail-call rule, the fresh label $\mathrm{tailCall}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In the application and tail-call cases, the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The six rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
@@ 4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces an elementary semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
@@
in the semantics of Section 3.
@@
Theorem 4.3 (First-order semantics of the lazy $\lambda$-calculus).
@@
Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2). The theorem establishes that the lazy $\lambda$-calculus has a first-order semantics: the translation $\mathcal{T}$ embeds the $\lambda$-calculus into a sublanguage of inheritance-calculus, using only recursive evaluation of set-comprehension equations over paths and sets of paths. The prior higher-order semantic frameworks of (Scott, 1972; Wadsworth, 1976; Abramsky et al., 2000; Hyland and Ong, 2000; Barendregt, 1984) are thereby shown to be faithful to this first-order semantics: they agree with the equations on Böhm tree equivalence (Barendregt, 1984). [^7] Composing $\mathcal{T}$ with the equations of Section 3 yields a first-order semantics of the $\lambda$-calculus in which function application is record inheritance and convergence is determined by path queries. Appendix E reformulates this semantics directly on ANF, without passing through inheritance-calculus. The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix B.
The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: contexts in the language universe $\mathscr{U}$ (Definition C.3) that mix $\lambda$-calculus and inheritance-calculus constructors can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathscr{U}}|_{L_{0}})}$ where $L_{0}$ is the lazy $\lambda$-calculus. This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem C.7).
@@ 4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent, and therefore indistinguishable in the $\lambda$-calculus, can be separated by a context that mixes $\lambda$-calculus and inheritance-calculus constructors, inheriting new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus (Theorem C.8): the $\lambda$-calculus can be macro-expressed in inheritance-calculus via ANF conversion and $\mathcal{T}$ (Lemma C.5, Theorem C.6), but the lazy $\lambda$-calculus cannot macro-express inheritance (Theorem C.7). The formal definitions, constructions, and proofs are given in Appendix C; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed: there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
@@ 5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names serve as data constructors and operations; lowerCamelCase names serve as fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
@@
Following the Church-encoding convention (Yang, 2026), [^8] The data representation (Zero/Successor) is a Peano encoding, an initial algebra of the functor $F(X)=1+X$. The visitor pattern (NatVisitor, below) adds the Church-encoding behavior: a fold (catamorphism) over this algebra. These two concerns are defined in separate files and composed via inheritance, which is itself an instance of the Expression Problem. we define a factory containing the abstract product type and two labels that serve as data constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
@@
Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each data constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
To test that $2+3=5$, we needed equality on Nat values. Equality requires case analysis on the data constructor; we use the visitor pattern. $\mathrm{NatVisitor}$ adds an $\mathrm{Acceptance}$ operation to each data constructor, containing a $\mathrm{VisitorMap}$ with the appropriate $\mathrm{ZeroVisitor}$ or $\mathrm{SuccessorVisitor}$ branch, and an $\mathrm{Accepted}$ field that delegates to the matching branch:
@@
$\quad\mathrm{Product}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{Accepted}\mapsto\{\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\}\},$
$\qquad\mathrm{Accepted}\mapsto\{\mathrm{VisitorMap}.\mathrm{ZeroVisitor}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{VisitorMap}\mapsto\{\mathrm{SuccessorVisitor}\mapsto\{\}\},$
$\qquad\mathrm{Accepted}\mapsto\{\mathrm{VisitorMap}.\mathrm{SuccessorVisitor}\}\}\}\}\}$
@@
$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each data constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Acceptance}$:
@@
$\hskip 18.49988pt\mathrm{OtherAcceptance}\mapsto\{\mathrm{other}.\mathrm{Acceptance},$
$\hskip 18.49988pt\;\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
$\hskip 18.49988pt\;\qquad\mathrm{SuccessorVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\}\},$
$\hskip 18.49988pt\;\mathrm{Accepted}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherAcceptance}.\mathrm{Accepted}.\mathrm{equal}\}\}\},$
@@
$\hskip 18.49988pt\mathrm{OtherAcceptance}\mapsto\{\mathrm{other}.\mathrm{Acceptance},$
$\hskip 18.49988pt\;\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\qquad\mathrm{SuccessorVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\}\},$
$\hskip 18.49988pt\;\mathrm{Accepted}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherAcceptance}.\mathrm{Accepted}.\mathrm{equal}\}\}\}\}\}$
@@
Each operation is an independent mixin that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every label in the factory acquires all operations from all inherited mixins. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential. In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix D), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
@@
A Church-encoded Nat is a record whose structure mirrors the data-constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$: